# HOPE: A Reinforcement Learning-based Hybrid Policy Path Planner for Diverse Parking Scenarios

### v1

@@ IV-C1 Efficient estimation of action mask
We first introduce a vectorized obstacle distance representation $l_{t}$ at timestamp t, where the $i^{th}$ element $l_{t}[i]$ is the nearest obstacle distance at angle $\omega_{i}=i\cdot\Delta\omega$ in the ego coordinate and $\Delta\omega$ is the angular resolution. Consider the envelope area $\text{S}_{\mathcal{E}}(\delta_{j},v\Delta t)$ covered by the vehicle traveling with a steering angle $\delta_{j}$ and step size $v\Delta t$. Let $\mathcal{E}_{ij}(v)$ denote the distance from the envelope boundary to the origin in the ego vehicle’s coordinate system at the $i^{th}$ angle $\omega_{i}$:
@@
The collision constraint can then be expressed as $\mathcal{E}_{ij}(v)\leq l_{t}[i]$. Then, the action mask calculation is equivalent to the following problem by introducing a variable $l$:
@@
While solving for the optimal $l_{ij}$ is equivalent to obtaining the maximum velocity $v^{*}_{ij}=\mathcal{E}_{ij}^{-1}(l_{t}[i])$ considering only obstacle point at $i^{th}$ angle $\omega_{i}$ and vehicle steering angle $\delta_{j}$, in practice, calculating all envelope distances corresponding to each $i$ and $j$ in each interaction is computationally expensive, and the inverse of $\mathcal{E}$ may not always exist. To deal with this issue, we propose the pre-calculation of the anchoring distance $\hat{l}_{ij}$ for $K$ discretized velocities:
@@
We take $v^{*}_{j}=\hat{v}[k^{*}_{j}]$ as a conservative estimation of max step velocity, and all actions at discretized angle $\delta_{j}$ are calculated in a vectorized manner at once. Since the anchoring distance $\hat{l}_{ij}$ is independent of information of obstacles at timestamp $t$, a matrix of all anchoring distance $\hat{\mathcal{L}}$ can be pre-calculated before the training process starts. This means that in each interaction step, calculating the action mask only requires the comparison between the two matrices, as shown in Equation 13, which simplifies the calculation.
@@ V Experiment
To better train and evaluate our approach, we rank the difficulty for parking scenarios into normal, complex, and extreme levels based on existing standards ISO 20900 and GB/T 41630-2022 [37, 38]. Generally, a parking space is defined by two boundary parking vehicles (or boundary obstacles) positioned on the two sides aligned along the curb-side edge of the road, as shown in Figure 2. The boundaries define the parking space’s length $L_{park}$ and width $W_{park}$. Other obstacles should be placed at least $D_{obst}$ away from the parking spots. We denote the parking vehicle’s width as $W$, length as $L$, and the distance from the start to the target position as $D_{park}$. Table I shows the categorization based on the above parameters. Since parallel parking is more challenging than vertical parking in practice, we introduce an extreme difficulty level for it. Besides, we rank the scenarios that $D_{park}>15.0$ the complex scenario. Note that we do not specify the initial orientation or position of the vehicle, meaning the starting conditions for the vehicle can be any collision-free configuration, which adds to the difficulty and diversity of scenarios.
@@ V-B Implement details
We conducted experiments in Tactics2D, an open-source simulator for driving decision-making [40]. This simulator provides sensor simulations, including lidar and bird’s-eye view. In each episode, the simulator independently and randomly initializes the parallel or vertical parking scenarios of three difficulty levels generated by the simulator or from the DLP dataset. Before the episode terminates, the simulator continually updates the vehicle state, calculates the reward, and returns it to the agent after the agent chooses an action. We conducted 100,000 episodes of training and tested 2,000 trials in each scenario category. Details about hyperparameters for algorithm and simulation are listed in Table VII in the appendix.
@@ V-C1 Comaparion with baselines
Table II presents the planning success rates of these methods across different scene types and difficulty levels. We have separately listed the results for DLP scenarios since they originate from real-world parking environments. As shown in the table, a direct implementation of the RS method fails to provide a feasible path in most cases. Hybrid A* achieves a success rate of over 90% in both normal scenarios and complex vertical parking cases. However, as the scene difficulty increases, its planning success rate significantly declines. In the extreme parallel parking scenarios, the most narrow cases, its success rate is only 16.8%. The reinforcement learning baselines perform better than the rule-based RS method but fail to surpass the Hybrid A*. Their low success rates in several scenarios demonstrate that achieving good performance through overfitting in a few cases does not necessarily guarantee the trained agent’s generalization capability. In contrast, our proposed HOPE, no matter based on on-policy PPO or off-policy SAC, outperforms all baselines and reaches a success rate of over 99.4% in all normal scenarios and over 94% in all scenarios. Figure 3 shows the reward and success rate curves in the training process. By combining the RL agent and the RS policy, the proposed method significantly improves training efficiency and success rate over the naive RL method.
@@ V-C2 Further comparison with Hybrid A*
Figure 4: The visualization of the planning process and results of the Hybrid A* (a-1)-(e-1) and the proposed HOPE (a-2)-(e-2). In the normal parallel parking case shown in (a), both methods provide concise path planning results. In the vertical parking scenario shown in (b) and the normal dlp scenario in (d), although both methods succeed in planning, our approach yields more reasonable results. Meanwhile, in the narrow parallel parking scenario (c) and the scenario requiring parking with the front of the vehicle facing inward (e), the Hybrid A* algorithm fails to plan, while our approach succeeds.
As the Hybrid A* is a widely used method to this day, we further compare our approach with it through specific case studies. As shown in Figure 4, while both two methods can mask successful path planning in some cases, our method is capable of providing more concise and reasonable planning results, such as in Figure 4 (b) and (d). As the parking space in the scene becomes narrow, Hybrid A* struggles to explore a feasible path. In contrast, our method, through training, achieves the ability to maneuver within tight parking spaces and overcome local optima situations, as shown in Figure 4 (c) and (e).
@@ V-D2 Other proposed modules
More ablation experiment results are presented in Table V. In the absence of the action mask mechanism, the success rate declines by more than 20% in complex parallel parking scenarios and 30% in extreme scenarios. This significant drop in performance indicates that utilizing an action mask to influence and constrain the agent’s action is crucial for training effectiveness. Besides, replacing the transformer with a multi-layer perceptron leads to a decline of more than 10% in several scenarios, demonstrating the importance of multi-modal information fusion. Excluding the action mask or bird’s-eye view (BEV) images as inputs to the neural network also impacts the agent’s performance. Using the auto-encoder to pre-train the image encoder can enhance the training for the PPO-based agent but has a relatively minor impact on the SAC-based agent.
@@ VI Conclusion and Future Work
To better train and evaluate the proposed method, we introduce a difficulty ranking approach for parking scenarios. The difficulty ranking standard categorizes scenarios into normal, complex, and extreme based on obstacles and parking space conditions. We utilized a simulator and a real-world parking lot dataset to construct training and testing scenarios. Comparisons were made against the Hybrid A* algorithm and the naive PPO and SAC. Comprehensive experiments were designed to illustrate the effectiveness of the proposed method. The results indicate that our approach outperforms the Hybrid A* algorithm in planning success rates across all types of scenarios, particularly in complex and extreme situations. Compared to the typical approach of directly using reinforcement learning for training, our proposed method combines the advantages of both rule-based and learning-based methods, resulting in an effectively trained planner with generalization capabilities across different scenarios.
This paper shows that the learning-based approach can serve as a promising tool in dealing with complex and diverse parking scenarios. Our work will be open-sourced to support the following research. We are planning to implement our algorithm within an automated parking system in an infrastructure-based indoor automated parking task. The proposed HOPE serves as a potential planner to deal with the intricate scenarios where the rule-based methods fail to plan a feasible path. Besides, since only static scenarios are considered in our work, more efforts could be made to apply the learning-based method in parking scenarios with dynamic obstacles in the future.
@@ Appendix A Implementation of information fusion transformer
  • Obstacle distance(vector-based) $l_{t}$: the value of the nearest distance to obstacles at certain angles, as defined in section IV-C1, which is similar to the scan results of a single-line lidar.
@@
  • Action mask(vector-based) $f_{am}$: a discretization of action mask that describes the max valid step size at different steering angles, as defined in section IV-C2.
  • Bird-eye-view information(image-based) $I_{BEV}$: a low-resolution depiction of the surrounding obstacle, target parking spot. The historical trajectory of the ego vehicle is rendered using a gradient color scheme. The image size is $H_{img}\cdot W_{img}$.
@@
Appendix B Reward function
@@
B-1 Intersection-of-union(IOU) reward
@@ B-2 Distance reward
B-3 Time consumption penalty
@@
B-4 Overall reward function

### v2

@@ IV-C1 Efficient estimation of action mask
We first introduce a vectorized obstacle distance representation $l_{t}$ at timestamp t, where the $i^{th}$ element $l_{t}[i]$ is the nearest obstacle distance at angle $\omega_{i}=i\cdot\Delta\omega$ in the ego coordinate and $\Delta\omega$ is the angular resolution. The movement distance during a certain period $\Delta t$ with some speed $v$ is denoted as step size $v\Delta t$. Consider the envelope area covered by the vehicle traveling with a steering angle $\delta_{j}$ and step size $v\Delta t$, denoted as $\text{S}_{\mathcal{E}}(\delta_{j},v\Delta t)$. Let $\mathcal{E}_{ij}(v)$ denote the distance from the envelope boundary to the origin in the ego vehicle’s coordinate system at the $i^{th}$ angle $\omega_{i}$:
@@
The collision constraint can then be expressed as $\mathcal{E}_{ij}(v)\leq l_{t}[i]$. Then, the action mask calculation is equivalent to the following problem by introducing a new 2-dim variable $l$:
@@
While solving for the optimal $l_{ij}$ considering only obstacle point at $i^{th}$ angle $\omega_{i}$ and vehicle steering angle $\delta_{j}$ is equivalent to obtaining the maximum velocity, denoted as $v^{*}_{ij}=\mathcal{E}_{ij}^{-1}(l_{t}[i])$, calculating all envelope distances corresponding to each $i$ and $j$ in each interaction is computationally expensive, and the inverse of $\mathcal{E}_{ij}$ may not always exist. To deal with this issue, we propose the pre-calculation of the anchoring distance $\hat{l}_{ij}$ for $K$ discretized velocities:
@@
We take $v^{*}_{j}=\hat{v}[k^{*}_{j}]$ as a conservative estimation of max step velocity, and all actions at discretized angle $\delta_{j}$ are calculated in a vectorized manner at once. Since the anchoring distance $\hat{l}_{ij}$ is independent of information of obstacles at timestamp $t$, a matrix of all anchoring distance $\hat{\mathcal{L}}$ can be pre-calculated before the training process starts. This implies that, as shown in Equation 13, calculating the action mask at each interaction step is simplified to a comparison between two matrices.
@@ V-A Scenario difficulty ranking
To better train and evaluate our approach, we rank the difficulty for parking scenarios into normal, complex, and extreme levels based on existing standards ISO 20900 and GB/T 41630-2022 [37, 38]. Generally, a parking space is defined by two boundary parking vehicles (or boundary obstacles) positioned on the two sides aligned along the curb-side edge of the road, as shown in Figure 2. The boundaries define the parking space’s length $L_{park}$ and width $W_{park}$. In our setup, other obstacles are located at least $D_{obst}$ away from the parking spots to leave the drivable space. We denote the parking vehicle’s width as $W$, length as $L$, and the distance from the start to the target position as $D_{park}$. Table I shows the categorization based on the aforementioned parameters, with narrower parking lots classified as more difficult. Since parallel parking is more challenging than vertical parking in practice, we introduce an extreme difficulty level for it [38]. Besides, because larger parking distances require more maneuvering over a longer distance and increase the number of obstacles encountered along the way, we rank the scenarios that $D_{park}>15.0$ the complex scenario. Note that we do not specify the initial orientation or position of the vehicle, meaning the starting conditions for the vehicle can be any collision-free configuration, which adds to the difficulty and diversity of scenarios.
@@ V-B Implement details
We conducted experiments in Tactics2D, an open-source simulator for driving decision-making [40]. This simulator provides sensor simulations, including lidar and bird’s-eye view (BEV). In each episode, the simulator independently and randomly initializes the parallel or vertical parking scenarios of three difficulty levels generated by the simulator or from the DLP dataset. We conducted 100,000 episodes of training in total and tested 2,000 trials in each scenario category, with no overlap between training and testing scenarios. Each time the scenario and its parameters are randomly generated according to the ranking method specified in section V-A Details about hyperparameters for algorithm and simulation are listed in Table VII in the appendix.
@@ V-C1 Comaparion with baselines
  • Experience-based heuristic search method (EBHS) [29]: The EBHS leverages reinforcement learning to train a Q-network that serves as a heuristic function within the heuristic search algorithm. This learning-based approach allows the algorithm to derive more suitable heuristic functions from data and improve the exploring efficiency.
@@
Table II presents the planning success rates of these methods across different scene types and difficulty levels. We have separately listed the results for DLP scenarios since they originate from real-world parking environments. As shown in the table, a direct implementation of the RS method fails to provide a feasible path in most cases. Hybrid A* achieves a success rate of over 90% in both normal scenarios and complex vertical parking cases. However, as the scene difficulty increases, its planning success rate significantly declines. In the extreme parallel parking scenarios, the most narrow cases, its success rate is only 16.8%. EBHS shows a significantly higher success rate in parallel parking scenarios than Hybrid A*. Specifically, in complex and extreme parallel parking scenarios, EBHS improves the success rate by nearly 30% over Hybrid A*. This improvement is attributed to the learning-based Q-function in EBHS, which can better estimate the state values in these scenarios compared to the rule-based heuristic function in Hybrid A*. However, the overall success rate of EBHS is still not high enough to be considered robust. The reinforcement learning baselines perform better than the rule-based RS method but fail to surpass the Hybrid A*. Their low success rates in several scenarios demonstrate that achieving good performance through overfitting in a few cases does not necessarily guarantee the trained agent’s generalization capability. In contrast, our proposed HOPE, no matter based on on-policy PPO or off-policy SAC, outperforms all baselines and reaches a success rate of over 99.4% in all normal scenarios and over 94% in all scenarios. Figure 3 shows the reward and success rate curves in the training process. The proposed method significantly improves training efficiency and success rate over the naive RL method by combining the RL agent and the RS policy.
@@ V-C2 Further comparison with Hybrid A*
Figure 4: The visualization of the planning process and results of the Hybrid A* (a-1)-(e-1) and the proposed HOPE (a-2)-(e-2). The blue-to-green gradient rectangles represent the states explored during the algorithm’s search process, while the yellow curves indicate the final path planning result. In the normal parallel parking case shown in (a), both methods provide concise path planning results. In the vertical parking scenario shown in (b) and the normal dlp scenario in (d), although both methods succeed in planning, our approach yields more reasonable results. Meanwhile, in the narrow parallel parking scenario (c) and the scenario requiring parking with the front of the vehicle facing inward (e), the Hybrid A* algorithm fails to plan, while our approach succeeds.
As the Hybrid A* is a widely used method to this day, we further compare our approach with it through specific case studies. As shown in Figure 4, while both two methods can mask successful path planning in some cases, our method is capable of providing more concise and reasonable planning results, such as in Figure 4 (b) and (d). As the parking space in the scene becomes narrow, Hybrid A* fails to explore a feasible path, as shown in Figure 4 (c) and (e). In contrast, our method, through training, achieves the ability to maneuver within tight parking spaces and overcome local optima situations
@@ V-D2 Other proposed modules
More ablation experiment results are presented in Table V. In the absence of the action mask mechanism, the success rate declines by more than 20% in complex parallel parking scenarios and 30% in extreme scenarios. This significant drop in performance indicates that utilizing an action mask to influence and constrain the agent’s action is crucial for training effectiveness. Besides, replacing the transformer with a multi-layer perceptron leads to a decline of more than 10% in several scenarios, demonstrating the importance of multi-modal information fusion. Excluding the action mask or BEV images as inputs to the neural network also impacts the agent’s performance. Using the auto-encoder to pre-train the image encoder can enhance the training for the PPO-based agent but has a relatively minor impact on the SAC-based agent.
@@ V-D3 Scenario difficulty
V-E Real-World Experiments
We also conducted real-world experiments in an indoor parking garage to show the applicability of our method in the real world. Our experimental platform was a Changan CS55E-Rock vehicle with a dimension of $4.6m\times 1.9m$ and we fine-tuned our model with the vehicle’s parameter in the same simulation environment. More details about the experimental platform and the overall system can be found in the appendix E.
We tested our method on vertical parking, parallel parking, and an additional challenging dead-end vertical parking scenario, as shown in Figure 6. To test the generalization ability of our algorithm, these scenarios are not included in the training set. In all three scenarios, our algorithm successfully planned collision-free paths and guided the vehicle to complete the parking process. Detailed videos of the parking processes can be found at https://www.youtube.com/watch?v=62w9qhjIuRI. It is worth noting that the dead-end scenario in Figure 6 (c) represents a situation where the drivable area in front of the parking spot is narrow, with one side obstructed by obstacles. This requires the vehicle to perform extensive maneuvering and multiple gear shifts to adjust its orientation, making it a highly challenging scenario. These results demonstrate the practical applicability and generalization capability of our algorithm in real-world scenarios.
((a))
((b))
((c))
((d))
((e))
((f))
Figure 6: The illustration of our real-world experiments in scenarios of vertical parking (a-1,2), parallel parking (b-1,2), and deadend parking (c-1,2). The initial position is shown in the left images and the right images show the target parking lots where the vehicle is positioned.
@@ VI Conclusion and Future Work
To better train and evaluate the proposed method, we introduce a difficulty ranking approach for parking scenarios. The difficulty ranking standard categorizes scenarios into normal, complex, and extreme based on obstacles and parking space conditions. We utilized a simulator and a real-world parking lot dataset to construct training and testing scenarios. Comparisons were made against other rule-based and RL-based methods. Comprehensive experiments were designed to illustrate the effectiveness of the proposed method. The results indicate that our approach outperforms the Hybrid A* algorithm in planning success rates across all types of scenarios, particularly in complex and extreme situations. Compared to the typical approach of directly using reinforcement learning for training, our proposed method combines the advantages of both rule-based and learning-based methods, resulting in an effectively trained planner with generalization capabilities across different scenarios.
This paper shows that the learning-based approach can serve as a promising tool in dealing with complex and diverse parking scenarios. Our work will be open-sourced to support the following research. The proposed HOPE serves as a potential planner to deal with the intricate scenarios where the rule-based methods fail to plan a feasible path. Besides, since only static scenarios are considered in our work, our method can be integrated with a dynamic obstacle avoidance system to handle real-world participants to ensure compatibility and safety in dynamic environments. In the future, more efforts could be made to apply the learning-based method directly in parking scenarios with dynamic obstacles. To support this, enhancing simulation environments with interactive traffic participant models that realistically react to the autonomous vehicle would greatly improve the training and evaluation of such methods.
@@ Appendix A Implementation of information fusion transformer
  • Obstacle distance (vector-based) $l_{t}$: the value of the nearest distance to obstacles at certain angles, as defined in section IV-C1, which is similar to the scan results of a single-line lidar. The size of this vector is fixed at 120 in practice, with each element representing the minimal obstacle distance within an angular range of $\pi/60$ radians.
@@
  • Action mask (vector-based) $f_{am}$: a discretization of action mask that describes the max valid step size at different steering angles, as defined in section IV-C2. The size of this vector is 42, with 21 elements representing the mask values for forward planning and the remaining 21 for backward planning.
  • Bird-eye-view information (image-based) $I_{BEV}$: a low-resolution depiction of the drivable area and target parking spot. The historical trajectory of the ego vehicle is rendered using a gradient color scheme. The image size is $H_{img}\cdot W_{img}$.
Figure 7: The success frequency of different RS curve types
@@
Appendix B Experiments on the Extended Reeds-Shepp Curves
We conducted experiments to evaluate the impact of different Reeds-Shepp (RS) curve types on the success rate of parking maneuvers. The results are summarized in Figure 7, where the horizontal axis represents different RS curve types, and the vertical axis represents the frequency of usage of each curve type in successful planning cases for all vertical parking scenarios. The results show that SCS curves have the second-highest frequency of 13.8%, demonstrating their effectiveness.
Appendix C Reward function
@@
C-1 Intersection-of-union(IOU) reward
@@ C-2 Distance reward
C-3 Time consumption penalty
@@
C-4 Overall reward function
@@ Appendix D Hyperparameters
$dim(l_{t})$ The dimension of obstacle distance vector $120$
$dim(f_{am})$ The dimension of action mask $42$
@@
Appendix E Real-world experiment setups
For sensors, we used Hesai Pandar40, RoboSense RS-LiDAR-32 and Bpearl LiDAR to support the vehicle’s localization and perception modules and no GPS was included. Figure 8 illustrates the framework of the parking system and how our method operates in real-world scenarios. The input to the planning module, where our method is applied, consists of a BEV grid map, obstacle distance, and the position of the parking spot. The target parking spot is selected by the driver and its precise location is obtained from a high-definition map, then transformed into the ego coordinate system through the localization system. Obstacle distance information is derived from the LiDAR point cloud, while the BEV grid map is generated by augmenting the drivable area from the high-definition map with real-time detected obstacles from the LiDAR. To mitigate the impact of upstream perception errors, we also add virtual occupancy grids in other parking spaces in the BEV grid map based on high-definition map information, ensuring the vehicle does not drive into these areas during maneuvering. These inputs are then converted into the format required by the network, which outputs a one-step planned path. Finally, a Quadratic Programming (QP) and Model Predictive Control (MPC) approach is used for path following to obtain the vehicle control commands from the waypoints. Besides, we also utilize a collision avoidance module to give way to dynamic obstacles in practice. All the perception and localization modules along with our planning algorithm are deployed on an Intel i7-1165G7 2.80GHz CPU.
Figure 8: The overall structure of the parking system, where our proposed planning module is in red.
Appendix F Proof of Convergence for the Hybrid Policy
The convergency of the update on the parameter $V_{\psi}$ and $Q_{\phi}$ can be proven by showing the update process of the Bellman expectation operator is a contraction mapping, which is independent of specific policy $\pi$. Thus the convergency stands for any given $\pi$ and so is the $\pi_{h}$. We here provide the proof of convergency of the value function $V_{\psi}$, and the Q function $Q_{\phi}$ can be obtained using a one-step iteration:
$Q^{\pi}(s,a)=\mathbb{E}\left[R_{t+1}+\gamma V^{\pi}(S_{t+1})\mid S_{t}=s,A_{t}=a\right]$ (22)
Firstly, the Bellman expectation equation for a policy $\pi$ is defined as:
$V_{\pi}(s_{t})=\mathbb{E}_{a_{t}\sim\pi(\cdot|s_{t})}[r(s_{t},a_{t})+\gamma V_{\pi}(s_{t+1})]$ (23)
We define the Bellman expectation operator $B_{\pi}$ as:
$(B_{\pi}V)(s_{t})=\mathbb{E}_{a\sim\pi(\cdot|s)}[r(s_{t},a)+\gamma V(s_{t+1})]$ (24)
To prove that the Bellman expectation operator $B_{\pi}$ is a contraction mapping, we need to show:
$\|B_{\pi}V_{1}-B_{\pi}V_{2}\|\leq\gamma\|V_{1}-V_{2}\|$ (25)
For any two value functions $V_{1}$ and $V_{2}$, we have:
$|B_{\pi}V_{1}(s_{t})-B_{\pi}V_{2}(s_{t})|$
$=\left|\mathbb{E}_{a\sim\pi(\cdot|s_{t})}\left[r(s_{t},a)+\gamma\mathbb{E}_{s_{t+1}\sim E}[V_{1}(s_{t+1})]\right]\right.$
$-\left.\mathbb{E}_{a\sim\pi(\cdot|s)}\left[r(s_{t},a)+\gamma\mathbb{E}_{s_{t+1}\sim E}[V_{2}(s_{t+1})]\right]\right|$
$=\gamma\left|\mathbb{E}_{a\sim\pi(\cdot|s_{t})}\left[\mathbb{E}_{s_{t+1}\sim E}[V_{1}(s_{t+1})-V_{2}(s_{t+1})]\right]\right|$
$=\gamma\left|\mathbb{E}_{a\sim\pi(\cdot|s_{t}),s_{t+1}\sim E}[V_{1}(s_{t+1})-V_{2}(s_{t+1})]\right|$
$\leq\gamma\mathbb{E}_{a\sim\pi(\cdot|s_{t}),s_{t+1}\sim P(\cdot|s_{t},a)}\left[\left|V_{1}(s_{t+1})-V_{2}(s_{t+1})\right|\right]$
$\leq\gamma\sup_{s_{t+1}}|V_{1}(s_{t+1})-V_{2}(s_{t+1})|$
$=\gamma\|V_{1}-V_{2}\|_{\infty}$
This shows that $B_{\pi}$ is a contraction mapping because $\|B_{\pi}V_{1}-B_{\pi}V_{2}\|_{\infty}\leq\gamma\|V_{1}-V_{2}\|_{\infty}$ for $10\leq\gamma<$ 1.
According to the Banach fixed-point theorem, a contraction mapping $B_{\pi}$ on a complete metric space $(\mathbb{R}^{S},\|\cdot\|_{\infty})$ has a unique fixed point $V_{\pi}^{*}$, i.e.,
$B_{\pi}V_{\pi}^{*}=V_{\pi}^{*}$ (26)
Therefore, through the iterative update process:
$V_{k+1}=B_{\pi}V_{k}$ (27)
the value function $V_{k}$ will converge.
The convergence proof above shows that the Bellman expectation operator $B_{\pi}$ is a contraction mapping and hence will converge to a fixed point for any given policy $\pi$. This means that the convergence of the value function updates is not dependent on the specific policy $\pi$. Therefore, even if we replace the policy $\pi$ with another policy $\pi_{h}$, the convergence properties still hold. We can use the following loss function to update the value function:
$J_{V}(\psi)=\mathbb{E}_{s_{t}\sim D_{\pi_{h}}}\left[\frac{1}{2}\left(V_{\psi}(s_{t})-\mathbb{E}_{\tilde{a}_{t}\sim\pi_{h}}\left[Q_{\phi}(s_{t},\tilde{a}_{t})\right]\right)^{2}\right]$ (28)
We minimize this loss function to update the parameter $\psi$:
$\psi_{k+1}=\psi_{k}-\alpha\nabla_{\psi}J_{V}(\psi_{k})$ (29)
This loss function minimizes the mean squared error between $V_{\psi}(s_{t})$ and $\mathbb{E}_{\tilde{a}_{t}\sim\pi_{h}}\left[Q_{\phi}(s_{t},\tilde{a}_{t})\right]$. According to the Bellman expectation equation:
$V_{\pi}(s_{t})=\mathbb{E}_{\tilde{a}_{t}\sim\pi_{h}}\left[Q_{\phi}(s_{t},\tilde{a}_{t})\right]$ (30)
minimizing $J_{V}(\psi)$ is equivalent to approximating the Bellman expectation equation.